# Efficient Parameter Optimisation for Quantum Kernel Alignment: A Sub-sampling Approach in Variational Training

### v1

@@ III.1.2 Learning Problem Labeling Cosets with Error (LCE)
Similarly to Section 2.1.1, we present the results obtained using the SPSA optimiser in Table 2 (see Appendix D Tables for the full set of results). Also in this case, we observe over a ×$\times 10$ speed-up, with the number of queries reduced as we reduce $k$. The classification accuracy of the method is, once again, not affected by the method and agrees with our previous discussion in Section 2.1.1.
@@
III.2 Real-World Dataset Experiment: Computational Pathology Breast Cancer Dataset
@@
Figure 2: Combined GNN-extracted Pati et al. [2022] hierarchical cellular features with our sub-sampling method to optimise the creation of a quantum kernel used to classify subtypes of cancer. The AI-extracted embeddings are used to create the dataset that will be sampled using the methodology discussed in this work. The output of the workflow is a fully variationally optimised quantum kernel.
@@
Accurate cancer detection and subtyping (i.e. the process of categorising a cancer into a more specific group based on its characteristics) O’Leary [2022], involves analyzing complex data from diagnostic images using Graph Neural Networks (GNNs). Typically the GNN is used to extract features that are then passed to a classifier Ahmedt-Aristizabal et al. [2022] (see Fig. 2). These classifiers have been shown to have high accuracy in classifying invasive vs non-invasive cancers but lower accuracy when classifying subtypes Pati et al. [2022].
In this work, we take the features extracted from a GNN using the framework as described in Pati et al. [2022] on breast cancer (BRACS) data. In the dataset, there are three distinct pathological benign lesions including Usual Ductal Hyperplasia (UDH), Flat Epithelial Atypia (FEA) and Atypical Ductal Hyperplasia (ADH), and malignant lesions including Ductal Carcinoma In Situ (DCIS) and Invasive Carcinoma (IC) besides Normal breast tissue. The subtypes are divided into two complex categories. The first category compares ADH and FEA to DCIS (1566 training points, 275 test points). In the second category Normal, and UDH are compared to ADH, FEA and DCIS (2797 Training points, 524 Test points).
In our tests, we used a 10-qubit feature map inspired by an instantaneous quantum polynomial (IQP) kernel Havlíček et al. [2019], Huang et al. [2021] with an optimal bandwidth parameter $c_{opt}\approx 2/n$ Shaydulin and Wild [2022], Canatar et al. [2022].
@@
The results highlighted in Fig. 3 are collected to show the performance of the sub-sampling method comparing all the optimisers at once. In the same Figure, we also make the comparison between the classification accuracy of a GNN, classical kernels and a standard quantum fidelity kernel. In the case of the sub-kernel method applied to the real-world problem of cancer subtyping, we observe similar trends to those observed for the synthetic datasets discussed above. The sub-sampling method does not degrade the classification performance, and in some cases it improves the F1 score accuracy by a small amount (top panels in Fig. 3). It is also evident that the quantum kernel with QKA outperforms the standard quantum fidelity kernel in terms of classification accuracy (blue dotted line in Figure 3).
The advantages of employing the sub-sampling method to accelerate kernel training are also clear (bottom panels in Fig. 3). Given the substantial size of the full kernels in these cases (ranging from $\sim$1600 to $\sim$2800 data points), the reduction in queries is even more significant than previously observed, with some instances requiring three orders of magnitude fewer queries. When a kernel of size $m=2796$ is reduced to a sub-kernel of size $k=8$ and $s=1$, the number of queries is reduced by a factor of $1.6\times 10^{3}$. A table with the full results is available in Appendix D Tables 10 - XIII.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption AF vs D Refer to caption NBU vs AFD
Figure 3: Top left and right panels, F1 scores obtained with our method, grouped by optimiser (ADAM, GD and SPSA), using the full dataset and for the classification of AF vs D and NBU vs AFD cancer subtypes. Quantum fidelity kernel, best classical kernel and GNN reference values are represented as horizontal lines. Bottom left and right panels: total number of queries for each method, grouped by optimiser.
IV Conclusion
@@
Acknowledgements.
This work was supported by the Hartree National Centre for Digital Innovation, a collaboration between the Science and Technology Facilities Council and IBM. We thank Stefan Woerner, Francesco Tacchino and Ivano Tavernelli for their valuable inputs. The project concept and research proposal was created and developed by Imperial College London and the Royal Brompton and Harefield Hospitals in collaboration with IBM. We also thank Voica Ana Marai Radescu, Michele Grossi, Chris Burton, Frederik Flöther (IBM) for their valuable support and input to the initial project development and concept.
@@ Appendix A Details of Experiments
Using a validation set, which is split from the training set, a grid-search is performed over $C$ and $\gamma$ parameters as well as kernel type (Poly, RFB, Linear), and the best parameters and kernel type are then used to find best Classical Kernel to train the SVC algorithm. Similarly, a grid-search is performed over different $C$ values for the obtained quantum kernels using the same validation set.
Appendix B Results for Hardware run of Synthetic Dataset with SPSA Optimizer
The hardware used in this work is IBM Quantum Nairobi with Falcon r5.11H processor and 7 qubits. For qubit connectivity refer to Fig. 4. Qiskit options were set as initial_layout=[1,2], transpile=3 and resillience-level=0 with 100 shots.
Figure 4: Schematic of the IBM Quantum Nairobi device taken at the time of the hardware experiments.
Table 3: Hardware results for SPSA Optimizer on the synthetic benchmarking dataset in section 2.1.1. $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze. The runtime is calculated as the total hardware execution time inside an IBM Quantum Session. Ansatz k s ROC AUC F1 Runtime (s) HE 96 1 0.85 0.721 20155 HE 32 1 0.904 0.906 2819 HE 32 8 0.94 0.889 16331 HE 16 4 0.995 0.953 3590 HE 4 8 0.992 0.953 2422 RA 96 1 0.86 0.80 19453 RA 32 1 0.807 0.694 2864 RA 32 8 0.859 0.766 15506 RA 16 4 0.891 0.812 3851
Appendix C Loss Graphs
The magnitudes of the loss values computed for different settings can vary considerably due to the use of different kernel sizes, leading to significantly different scales of loss. Direct comparisons of these values are not informative because of this scale variation. To facilitate meaningful comparisons, min-max scaling is applied to each loss graph in Figs. 5 - 7, standardizing the range of values. The purpose of these plots is to demonstrate that, as expected, the sub-sampling method does result in the loss converging at a slower rate compared to the full kernel method.
Refer to captionRefer to caption
@@ Appendix C Loss Graphs
Appendix D Numerical Results
@@
Table 4: Results for SPSA Optimizer on the synthetic benchmarking dataset in section 2.1.1. $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze.
Ansatz k s ROC AUC F1 Queries Speed-up
HE 96 1 0.9800 0.9312 1377696.00 1.00
HE 32 1 0.9900 0.9312 378848.00 3.64
HE 32 4 0.9850 0.9312 1380320.00 1.00
HE 32 8 0.9800 0.9312 3125216.00 0.44
HE 16 1 0.9850 0.8561 105712.00 13.03
HE 16 4 0.9900 0.9312 452080.00 3.05
HE 16 8 0.9850 0.9312 894960.00 1.54
HE 8 1 0.9800 0.9312 14808.00 93.04
HE 8 4 0.9850 0.9312 108536.00 12.69
HE 8 8 0.9850 0.9312 213752.00 6.45
HE 4 1 0.9600 0.7407 3012.00 457.40
HE 4 4 0.9650 0.8667 26908.00 51.20
HE 4 8 1.0000 0.9312 54460.00 25.30
RA 96 1 0.9800 0.9312 852384.00 1.00
RA 32 1 0.9850 0.9312 33760.00 25.25
RA 32 4 0.9850 0.9312 1558496.00 0.55
RA 32 8 0.9900 0.9312 3063776.00 0.28
RA 16 1 0.9850 0.9312 87024.00 9.79
RA 16 4 0.9850 0.9312 260784.00 3.26
RA 16 8 0.8900 0.8244 919536.00 0.93
RA 8 1 0.9800 0.8624 27960.00 30.49
RA 8 4 0.9700 0.8986 101496.00 8.40
RA 8 8 0.9750 0.8986 227064.00 3.75
RA 4 1 0.9050 0.8244 4972.00 171.44
RA 4 4 0.9700 0.8986 28636.00 29.77
RA 4 8 0.8650 0.7634 56700.00 15.03
@@
Table 5: Results for ADAM Optimizer on the synthetic benchmarking dataset in section 2.1.1. $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze.
Ansatz k s ROC AUC F1 Queries Speed-up
HE 96 1 0.9850 0.9312 6114720.00 1.00
HE 32 1 0.9700 0.7407 2157024.00 2.83
HE 32 4 0.9150 0.8244 3012576.00 2.03
HE 32 8 0.9600 0.8624 491488.00 12.44
HE 16 1 0.9500 0.8561 111472.00 54.85
HE 16 4 0.9550 0.8624 1893872.00 3.23
HE 16 8 0.9800 0.9312 902128.00 6.78
HE 8 1 0.9800 0.9312 344.00 17775.35
HE 8 4 0.9750 0.9662 689400.00 8.87
HE 8 8 0.9750 0.8624 158456.00 38.59
HE 4 1 0.9150 0.7937 60.00 101912.00
HE 4 4 0.9850 0.9312 148220.00 41.25
HE 4 8 0.9550 0.8624 344956.00 17.73
RA 96 1 0.9100 0.5333 350112.00 1.00
RA 32 1 0.9400 0.8986 56800.00 6.16
RA 32 4 0.9650 0.8244 88032.00 3.98
RA 32 8 0.9050 0.8000 9957344.00 0.04
RA 16 1 0.8700 0.6936 66416.00 5.27
RA 16 4 0.9850 0.8624 68080.00 5.14
RA 16 8 0.9750 0.8624 984048.00 0.36
RA 8 1 0.8500 0.8000 184.00 1902.78
RA 8 4 0.9150 0.8148 383736.00 0.91
RA 8 8 0.8700 0.7937 768248.00 0.46
RA 4 1 0.8950 0.8244 116.00 3018.21
RA 4 4 0.9700 0.9662 32060.00 10.92
RA 4 8 0.9300 0.8624 172732.00 2.03
@@
Table 6: Results for GD Optimizer on the synthetic benchmarking dataset in section 2.1.1. $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze.
Ansatz k s ROC AUC F1 Queries Speed-up
HE 96 1 0.9800 0.9312 6437280.00 1.00
HE 32 1 0.9800 0.9312 1163232.00 5.53
HE 32 4 0.9850 0.9312 6948832.00 0.93
HE 32 8 0.9800 0.9312 13897696.00 0.46
HE 16 1 0.9800 0.9312 323952.00 19.87
HE 16 4 0.9900 0.8947 2047984.00 3.14
HE 16 8 0.9800 0.9312 4088816.00 1.57
HE 8 1 0.9700 0.8624 2520.00 2554.48
HE 8 4 0.9650 0.8947 511992.00 12.57
HE 8 8 0.9650 0.8986 1023224.00 6.29
HE 4 1 0.9750 0.8624 31876.00 201.95
HE 4 4 0.9750 0.8561 68668.00 93.74
HE 4 8 0.8500 0.7542 189180.00 34.03
RA 96 1 0.9800 0.9312 2386848.00 1.00
RA 32 1 0.9150 0.8310 1036768.00 2.30
RA 32 4 0.9850 0.9312 4171744.00 0.57
RA 32 8 0.8500 0.7333 8347616.00 0.29
RA 16 1 0.8650 0.5333 2416.00 987.93
RA 16 4 0.9850 0.9312 1226736.00 1.95
RA 16 8 0.9900 0.9312 2455536.00 0.97
RA 8 1 0.9450 0.8667 2520.00 947.16
RA 8 4 0.9300 0.8310 192376.00 12.41
RA 8 8 0.9700 0.8624 614392.00 3.88
RA 4 1 0.8100 0.7249 6244.00 382.26
RA 4 4 0.9850 0.8624 76604.00 31.16
RA 4 8 0.9150 0.7634 13564.00 175.97
@@
Table 7: Results for SPSA Optimizer on the COSET dataset in section 2.1.2. $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze.
Ansatz k s ROC AUC F1 Queries Speed-up
HE 96 1 1.0000 1.0000 1396128.00 1.00
HE 32 1 1.0000 1.0000 382944.00 3.65
HE 32 4 1.0000 1.0000 1542112.00 0.91
HE 32 8 1.0000 1.0000 3125216.00 0.45
HE 16 1 1.0000 1.0000 113136.00 12.34
HE 16 4 1.0000 1.0000 455152.00 3.07
HE 16 8 1.0000 0.9687 901104.00 1.55
HE 8 1 0.9569 0.8110 11736.00 118.96
HE 8 4 0.8510 0.7500 96504.00 14.47
HE 8 8 0.7765 0.7500 228600.00 6.11
HE 4 1 0.7412 0.7500 1668.00 837.01
HE 4 4 0.7137 0.7500 18940.00 73.71
HE 4 8 0.6941 0.7179 35452.00 39.38
RA 96 1 1.0000 1.0000 1428384.00 1.00
RA 32 1 1.0000 1.0000 385504.00 3.71
RA 32 4 1.0000 1.0000 1507296.00 0.95
RA 32 8 1.0000 1.0000 3084256.00 0.46
RA 16 1 1.0000 1.0000 106224.00 13.45
RA 16 4 1.0000 1.0000 446448.00 3.20
RA 16 8 1.0000 0.9687 917488.00 1.56
RA 8 1 0.6157 0.6161 22040.00 64.81
RA 8 4 0.9255 0.8110 115064.00 12.41
RA 8 8 0.9137 0.8110 229880.00 6.21
RA 4 1 0.7882 0.7500 2972.00 480.61
RA 4 4 0.6980 0.7179 27516.00 51.91
RA 4 8 0.7686 0.7500 33660.00 42.44
@@
Table 8: Results for ADAM Optimizer on the COSET dataset in section 2.1.2. $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze.
Ansatz k s ROC AUC F1 Queries Speed-up
HE 96 1 1.0000 1.0000 7741344.00 1.00
HE 32 1 0.9137 0.8110 2088928.00 3.71
HE 32 4 0.9765 0.8110 8173536.00 0.95
HE 32 8 0.9804 0.9375 5533664.00 1.40
HE 16 1 0.9176 0.8125 614512.00 12.60
HE 16 4 0.7176 0.7179 1090544.00 7.10
HE 16 8 0.8118 0.7500 4916208.00 1.57
HE 8 1 1.0000 1.0000 153272.00 50.51
HE 8 4 0.6941 0.7179 18936.00 408.82
HE 8 8 0.7255 0.7500 146424.00 52.87
HE 4 1 0.9843 0.9060 38268.00 202.29
HE 4 4 0.9725 0.8740 153500.00 50.43
HE 4 8 1.0000 1.0000 306812.00 25.23
RA 96 1 1.0000 1.0000 7741344.00 1.00
RA 32 1 0.7647 0.7500 1883616.00 4.11
RA 32 4 0.8980 0.8110 8355808.00 0.93
RA 32 8 0.8627 0.7806 16711648.00 0.46
RA 16 1 0.9961 0.8414 613872.00 12.61
RA 16 4 0.8118 0.7500 2391024.00 3.24
RA 16 8 0.7216 0.7500 3724272.00 2.08
RA 8 1 1.0000 1.0000 153528.00 50.42
RA 8 4 0.7373 0.7500 613496.00 12.62
RA 8 8 0.8275 0.7500 1229048.00 6.30
RA 4 1 1.0000 1.0000 38388.00 201.66
RA 4 4 1.0000 1.0000 153468.00 50.44
RA 4 8 0.7294 0.7500 137980.00 56.10

### v2

@@ 3.1.2 Learning Problem Labeling Cosets with Error (LCE)
Similarly to Section 3.1.1, we present the results obtained using the SPSA optimiser in Figure 2 (see Appendix E Tables for the full set of results). Also in this case, we observe over a ×$\times 10$ speed-up, with the number of queries reduced as we reduce $k$. For both the HE and RA ansatzes in Fig. 2, the full kernel ROC AUC and F1 score are 1. For all choices of $k$ and $s$ the ROC AUC remains at 1. All F1 scores are also 1 apart from the choice of $k=16$, $s=8$ which sees a small reduction to 0.969. We therefore see that the classification accuracy of the method is, once again, not affected by the method and agrees with our previous discussion in Section 3.1.1.
@@
3.2 Real-World Dataset Experiment: Computational Pathology Breast Cancer Dataset
[…]
[width=1]Rebuttal-for-quantum/plots/schematic_pathology.pdf
@@ 3.2 Real-World Dataset Experiment: Computational Pathology Breast Cancer Dataset
Figure 3: Combined GNN-extracted [27] hierarchical cellular features with our sub-sampling method to optimise the creation of a quantum kernel used to classify subtypes of cancer. The GNN-extracted embeddings are used to create the dataset that will be sampled using the methodology discussed in this work. The output of the workflow is a fully variationally optimised quantum kernel.
@@
Accurate cancer detection and subtyping (i.e. the process of categorising a cancer into a more specific group based on its characteristics) [26], involves analyzing complex data from diagnostic images using Graph Neural Networks (GNNs). Typically the GNN is used to extract features that are then passed to a classifier [1] (see Fig. 3). These classifiers have been shown to have high accuracy in classifying invasive vs non-invasive cancers but lower accuracy when classifying subtypes [27].
In this work, we take the features extracted from a GNN using the framework as described in [27] on breast cancer (BRACS) data. In the dataset, there are three distinct pathological benign lesions including Usual Ductal Hyperplasia (UDH), Flat Epithelial Atypia (FEA) and Atypical Ductal Hyperplasia (ADH), and malignant lesions including Ductal Carcinoma In Situ (DCIS) and Invasive Carcinoma (IC) besides Normal breast tissue. The subtypes are divided into two complex categories. The first category compares ADH and FEA to DCIS (1566 training points, 275 test points). In the second category Normal, and UDH are compared to ADH, FEA and DCIS (2797 Training points, 524 Test points).
In our tests, we used a 10-qubit feature map inspired by an instantaneous quantum polynomial (IQP) kernel [15, 17] with an optimal bandwidth parameter $c_{opt}\approx 2/n$ [31, 6] to reduce effects of exponential concentration quantum kernels [34].
@@
The results highlighted in Fig. 4 are collected to show the performance of the sub-sampling method comparing all the optimisers at once. In the same Figure, we also make the comparison between the classification accuracy of a GNN, classical kernels, PEGASOS kernel alignment [12] and a quantum fidelity kernel without kernel alignment. In the case of the sub-kernel method applied to the real-world problem of cancer subtyping, we observe similar trends to those observed for the synthetic datasets discussed above. The sub-sampling method does not degrade the classification performance,(top panels in Fig. 4). It is also evident that the quantum kernel with QKA outperforms both the quantum fidelity kernel without kernel alignment and PEGASOS kernel alignment in terms of classification accuracy (cyan line and blue dotted line in Fig. 4). The comparison of the results for the breast cancer data set between the sub-kernel method and the classical GNN approach indicates that the sub-kernel method can achieve comparable results to the classical method. While the initial findings are promising, they should be interpreted as indicative rather than conclusive.
The advantages of employing the sub-sampling method to accelerate kernel training are also clear (bottom panels in Fig. 4). Given the substantial size of the full kernels in these cases (ranging from $\sim$1600 to $\sim$2800 data points), the reduction in queries is even more significant than previously observed, with some instances requiring three orders of magnitude fewer queries. When a kernel of size $m=2796$ is reduced to a sub-kernel of size $k=8$ and $s=1$, the number of queries is reduced by a factor of $1.6\times 10^{3}$. These results provide some evidence that it is possible to choose the parameters $s$ and $k$ as constants independent of $m$. In the case that $m=1566$, we find that choosing $k=16$ and $s=1$ gives a worst case reduction in ROC AUC and F1 score of 1.77% and 1.02% respectively. On average ROC AUC increases by 1.16% and the F1 score decreases by 0.1% (see Appendix E Table X). Increasing the full kernel size to $m=2796$ gives similarly good results for the same parameters, $k=16$ and $s=1$. The worst case reduction in ROC and F1 score is 1.18% and 0.36% respectively. On average the ROC AUC increases by 0.61% the F1 score increases by 0.02% (see Appendix E Table XII). While we can’t generalise from these two data points, the fact that we observe that $k$ and $s$ do not need to grow with $m$ in this case is an encouraging sign that the sub-kernel method can provide highly efficient scaling, potentially independent of $m$.
[…]
Figure 4: Top left and right panels, F1 scores obtained with our method, grouped by optimiser (ADAM, GD and SPSA), using the full dataset and for the classification of AF vs D and NBU vs AFD cancer subtypes. Quantum fidelity kernel, best classical kernel and GNN reference values are represented as horizontal lines. Bottom left and right panels: total number of queries for each method, grouped by optimiser. The best results are obtained after a parameter grid search as described in A, and full results for PEGASOS QKA is given in B
4 Conclusion
@@
This work was supported by the Hartree National Centre for Digital Innovation, a collaboration between the Science and Technology Facilities Council and IBM. We thank Stefan Woerner, Francesco Tacchino and Ivano Tavernelli for their valuable inputs. The project concept and research proposal was created and developed by Imperial College London and the Royal Brompton and Harefield Hospitals in collaboration with IBM. We also thank Voica Ana Maria Radescu, Michele Grossi, Chris Burton, Frederik Flöther for their valuable support and input to the initial project development and concept.
@@ Appendix A Details of Experiments
Using a validation set, which is split from the training set, a grid-search is performed over $C$ and $\gamma$ parameters as well as kernel type (Poly, RFB, Linear), and the best parameters and kernel type are then used to find best Classical Kernel to train the SVC algorithm. Similarly, a grid-search is performed over different $C$ values for the obtained quantum kernels using the same validation set for both QKA and PEGASOS QKA algorithms.
Appendix B Results of PEGASOS QKA
Table 1: Reference Values for NBU vs AFD and AF vs D. Best results of PEGASOS QKA [12] from the grid search as described in Appendix A
NBU vs AFD
Number of Steps F1 Score
10 0.602
100 0.706
1,000 0.765
10,000 0.787
100,000 0.780
AF vs D
Number of Steps F1 Score
10 0.737
100 0.766
1,000 0.763
10,000 0.760
100,000 0.760
Appendix C Results for Hardware run of Synthetic Dataset with SPSA Optimizer
The hardware used in this work is IBM’s ibm_nairobi with Falcon r5.11H processor and 7 qubits. For qubit connectivity refer to Fig. 5. Qiskit options were set as initial_layout=[1,2], transpile=3 and resillience-level=0 with 100 shots.
[…]
[width=0.2]Rebuttal-for-quantum/Schematics/nairobi.png
Figure 5: Schematic of the IBM quantum hardware ibm_nairobi device taken at the time of the hardware experiments.
[…]
! Table 2: It has been only run once. Hardware results for SPSA Optimizer on the synthetic benchmarking dataset in section 2.1.1. $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze. The runtime is calculated as the total hardware execution time inside an IBM Quantum Session. Ansatz k s ROC AUC F1 Runtime (s) HE 96 1 0.85 0.721 20155 HE 32 1 0.904 0.906 2819 HE 32 8 0.94 0.889 16331 HE 16 4 0.995 0.953 3590 HE 4 8 0.992 0.953 2422 RA 96 1 0.86 0.8 19453 RA 32 1 0.807 0.694 2864 RA 32 8 0.859 0.766 15506 RA 16 4 0.891 0.812 3851
Appendix D Loss Graphs
The magnitudes of the loss values computed for different settings can vary considerably due to the use of different kernel sizes, leading to significantly different scales of loss. Direct comparisons of these values are not informative because of this scale variation. To facilitate meaningful comparisons, min-max scaling is applied to each loss graph in Figs. 6 - 8, standardizing the range of values. The purpose of these plots is to demonstrate that, as expected, the sub-sampling method does result in the loss converging at a slower rate compared to the full kernel method.
[…]
@@ Appendix D Loss Graphs
Appendix E Numerical Results
@@
Table 3: Results for SPSA Optimizer on the synthetic benchmarking dataset in section 2.1.1. $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze. Full kernels ($k=m$) are highlighted in bold.
[…]
0.4! Ansatz k s ROC AUC F1 Queries Speed-up HE 96 1 0.98 0.9312 1377696 1 HE 32 1 0.99 0.9312 378848 3.64 HE 32 4 0.985 0.9312 1380320 1 HE 32 8 0.98 0.9312 3125216 0.44 HE 16 1 0.985 0.8561 105712 13.03 HE 16 4 0.99 0.9312 452080 3.05 HE 16 8 0.985 0.9312 894960 1.54 HE 8 1 0.98 0.9312 14808 93.04 HE 8 4 0.985 0.9312 108536 12.69 HE 8 8 0.985 0.9312 213752 6.45 HE 4 1 0.96 0.7407 3012 457.40 HE 4 4 0.965 0.8667 26908 51.20 HE 4 8 1 0.9312 54460 25.30 RA 96 1 0.98 0.9312 852384 1 RA 32 1 0.985 0.9312 33760 25.25 RA 32 4 0.985 0.9312 1558496 0.55 RA 32 8 0.990 0.9312 3063776 0.28 RA 16 1 0.985 0.9312 87024 9.79 RA 16 4 0.985 0.9312 260784 3.26 RA 16 8 0.89 0.8244 919536 0.93 RA 8 1 0.98 0.8624 27960 30.49 RA 8 4 0.97 0.8986 101496 8.4 RA 8 8 0.975 0.8986 227064 3.75 RA 4 1 0.905 0.8244 4972 171.44 RA 4 4 0.97 0.8986 28636 29.77 RA 4 8 0.865 0.7634 56700 15.03
@@
Table 4: Results for ADAM Optimizer on the synthetic benchmarking dataset in section 2.1.1. $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze. Full kernels ($k=m$) are highlighted in bold.
[…]
0.4! Ansatz k s ROC AUC F1 Queries Speed-up HE 96 1 0.985 0.9312 6114720 1 HE 32 1 0.97 0.7407 2157024 2.83 HE 32 4 0.915 0.8244 3012576 2.03 HE 32 8 0.96 0.8624 491488 12.44 HE 16 1 0.95 0.8561 111472 54.85 HE 16 4 0.955 0.8624 1893872 3.23 HE 16 8 0.98 0.9312 902128 6.78 HE 8 1 0.98 0.9312 344 17775.35 HE 8 4 0.975 0.9662 689400 8.87 HE 8 8 0.975 0.8624 158456 38.59 HE 4 1 0.915 0.7937 60 101912 HE 4 4 0.985 0.9312 148220 41.25 HE 4 8 0.955 0.8624 344956 17.73 RA 96 1 0.91 0.5333 350112 1 RA 32 1 0.94 0.8986 56800 6.16 RA 32 4 0.965 0.8244 88032 3.98 RA 32 8 0.905 0.8 9957344 0.04 RA 16 1 0.87 0.6936 66416 5.27 RA 16 4 0.985 0.8624 68080 5.14 RA 16 8 0.975 0.8624 984048 0.36 RA 8 1 0.85 0.8000 184 1902.78 RA 8 4 0.915 0.8148 383736 0.91 RA 8 8 0.87 0.7937 768248 0.46 RA 4 1 0.895 0.8244 116 3018.21 RA 4 4 0.97 0.9662 32060 10.92 RA 4 8 0.93 0.8624 172732 2.03
@@
Table 5: Results for GD Optimizer on the synthetic benchmarking dataset in section 2.1.1. $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze. Full $k$ are highlighted in bold.
[…]
0.4! Ansatz k s ROC AUC F1 Queries Speed-up HE 96 1 0.98 0.9312 6437280 1 HE 32 1 0.98 0.9312 1163232 5.53 HE 32 4 0.985 0.9312 6948832 0.93 HE 32 8 0.98 0.9312 13897696 0.46 HE 16 1 0.98 0.9312 323952 19.87 HE 16 4 0.99 0.8947 2047984 3.14 HE 16 8 0.98 0.9312 4088816 1.57 HE 8 1 0.97 0.8624 2520 2554.48 HE 8 4 0.965 0.8947 511992 12.57 HE 8 8 0.965 0.8986 1023224 6.29 HE 4 1 0.975 0.8624 31876 201.95 HE 4 4 0.975 0.8561 68668 93.74 HE 4 8 0.85 0.7542 189180 34.03 RA 96 1 0.98 0.9312 2386848 1 RA 32 1 0.915 0.831 1036768 2.3 RA 32 4 0.985 0.9312 4171744 0.57 RA 32 8 0.85 0.7333 8347616 0.29 RA 16 1 0.865 0.5333 2416 987.93 RA 16 4 0.985 0.9312 1226736 1.95 RA 16 8 0.99 0.9312 2455536 0.97 RA 8 1 0.945 0.8667 2520 947.16 RA 8 4 0.93 0.831 192376 12.41 RA 8 8 0.97 0.8624 614392 3.88 RA 4 1 0.81 0.7249 6244 382.26 RA 4 4 0.985 0.8624 76604 31.16 RA 4 8 0.915 0.7634 13564 175.97
@@
Table 6: Results for SPSA Optimizer on the COSET dataset in section 2.1.2. $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze. Full kernels ($k=m$) are highlighted in bold.
[…]
0.4! Ansatz k s ROC AUC F1 Queries Speed-up HE 96 1 1 1 1396128 1 HE 32 1 1 1 382944 3.65 HE 32 4 1 1 1542112 0.91 HE 32 8 1 1 3125216 0.45 HE 16 1 1 1 113136 12.34 HE 16 4 1 1 455152 3.07 HE 16 8 1 0.9687 901104 1.55 HE 8 1 0.9569 0.811 11736 118.96 HE 8 4 0.8510 0.75 96504 14.47 HE 8 8 0.7765 0.75 228600 6.11 HE 4 1 0.7412 0.75 1668 837.01 HE 4 4 0.7137 0.75 18940 73.71 HE 4 8 0.6941 0.7179 35452 39.38 RA 96 1 1 1 1428384 1 RA 32 1 1 1 385504 3.71 RA 32 4 1 1 1507296 0.95 RA 32 8 1 1 3084256 0.46 RA 16 1 1 1 106224 13.45 RA 16 4 1 1 446448 3.20 RA 16 8 1 0.9687 917488 1.56 RA 8 1 0.6157 0.6161 22040 64.81 RA 8 4 0.9255 0.811 115064 12.41 RA 8 8 0.9137 0.811 229880 6.21 RA 4 1 0.7882 0.75 2972 480.61 RA 4 4 0.6980 0.7179 27516 51.91 RA 4 8 0.7686 0.75 33660 42.44
@@
Table 7: Results for ADAM Optimizer on the COSET dataset in section 2.1.2. $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze. Full kernels ($k=m$) are highlighted in bold.
[…]
0.4! Ansatz k s ROC AUC F1 Queries Speed-up HE 96 1 1 1 7741344 1 HE 32 1 0.9137 0.811 2088928 3.71 HE 32 4 0.9765 0.811 8173536 0.95 HE 32 8 0.9804 0.9375 5533664 1.40 HE 16 1 0.9176 0.8125 614512 12.60 HE 16 4 0.7176 0.7179 1090544 7.10 HE 16 8 0.8118 0.75 4916208 1.57 HE 8 1 1 1 153272 50.51 HE 8 4 0.6941 0.7179 18936 408.82 HE 8 8 0.7255 0.75 146424 52.87 HE 4 1 0.9843 0.906 38268 202.29 HE 4 4 0.9725 0.874 153500 50.43 HE 4 8 1 1 306812 25.23 RA 96 1 1 1 7741344 1 RA 32 1 0.7647 0.75 1883616 4.11 RA 32 4 0.8980 0.811 8355808 0.93 RA 32 8 0.8627 0.7806 16711648 0.46 RA 16 1 0.9961 0.8414 613872 12.61 RA 16 4 0.8118 0.75 2391024 3.24 RA 16 8 0.7216 0.75 3724272 2.08 RA 8 1 1 1 153528 50.42 RA 8 4 0.7373 0.75 613496 12.62 RA 8 8 0.8275 0.75 1229048 6.3 RA 4 1 1 1 38388 201.66 RA 4 4 1 1 153468 50.44 RA 4 8 0.7294 0.75 137980 56.1